OpenStudy (anonymous):

how do you fine rad with the calculator...3rad divided by 4...

OpenStudy (anonymous):

rad meaning radians. Its different for each calc unfortunately but you can work it out for yourself because deg = \[degrees \times \frac{\pi}{180} = radians\]

OpenStudy (anonymous):

there are 3.142 radians in PI which is 180 degrees

OpenStudy (anonymous):

oh ok so explain how 18 degrees . rad over 180 degrees = 0.3 rad

OpenStudy (anonymous):

3 radians has a value of \[3*\frac{180}{\pi}=171.887 degrees\] and i guess \[\frac{171.887}{4}=42.97175\]

OpenStudy (anonymous):

\[18 \times \frac{\pi}{180} = 0.314\]

OpenStudy (anonymous):

18 degrees is a 10th of 180 so it must be a tenth of pi

OpenStudy (anonymous):

If you times your answer by 10 you will get PI back. see?

OpenStudy (anonymous):

yes that makes sense...thks...